OpenStudy (anonymous):

if a snail teravels at 3 3/4 feet an hour how many feet does a snail traval in 3 1/2

OpenStudy (anonymous):

fan and medal

Mehek (mehek14):

3 1/2 hours?

OpenStudy (anonymous):

yes

Mehek (mehek14):

ok 3.75 per hour 3.75 * 3.5 multiply that

OpenStudy (anonymous):

11.025

Mehek (mehek14):

no...

Mehek (mehek14):

it is 13.125

OpenStudy (anonymous):

ok thanks can you stay on if I need more hlp plz

Mehek (mehek14):

wait do you need the answer in fraction form?

OpenStudy (anonymous):

yes

Mehek (mehek14):

ok so it would be 13 1/8

OpenStudy (anonymous):

Which arrangement shows , , , and in order from least to greatest? A. B. C. D.

OpenStudy (anonymous):

@Mehek14

Mehek (mehek14):

no numbers?

OpenStudy (anonymous):

27/5, 5.36, 5 3/5,and 33/6 sry

Mehek (mehek14):

change all to decimals 27/5 = 5.4 5.36 5 3/5 = 5.6 33/6 = 5.5 so 5.36, 27/5, 33/6, 5 3/5

OpenStudy (anonymous):

What is the perimeter of a rectangle with a length of 6 4/5 and a width of 2 2/3 @Mehek14

Mehek (mehek14):

34/5*2 + 8/3*2

OpenStudy (anonymous):

um can you tell me the answer please

Mehek (mehek14):

5.33 + 13.6 18.93

OpenStudy (anonymous):

whats that as a fraction

Mehek (mehek14):

ok now i have to gtg so i can't help with more now

Mehek (mehek14):

it's 18 93/100